# ARC: Argument Representation and Coverage Analysis for Zero-Shot Long Document Summarization with Instruction Following LLMs

### v1

@@ 4.3 Obtaining generated summaries
To handle the long-form nature of argumentative texts, we employ open-weight LLMs with extended context windows: LlaMA-3.1-8B-instruct (LlaMA-3.18-B) Grattafiori et al. (2024), Mistral-8B-instruct(Mistral-8b) Jiang et al. (2024), and Qwen2-7B qwe (2024). For Qwen2-7B, to support context lengths exceeding 32k tokens, we integrate YARN embeddings Peng et al. into the model configuration prior to deployment. Inference is conducted using VLLM Kwon et al. (2023) for scalability. Summaries are generated using temperature sampling, capped at tokens to ensure fair long-form generation. Each document is prompted with: "Read the following text and summarize it: {input document}. Summarize in {reference summary word length} words. Summary:". The target length is dynamically set to match reference summaries for comparability. For DRI, the length is fixed to the longest reference summary to encourage maximal argument retention.
5 Results and Analysis
5.1 Do LLMs cover salient arguments effectively?
We compute ARC888Full ARC and ARC results are in Appendix F. scores for human reference summaries on both the CANLII and DRI benchmarks using GPT-4o. For CANLII, this evaluation serves to assess the robustness of the metric by verifying whether it assigns near-perfect scores to ideal human-written summaries. For DRI, the goal is to measure how much human references cover salient argument roles compared to LLM-generated summaries. As shown in Figure 2, argument coverage remains imperfect across all models, with Mistral-8B lagging behind both LlaMA-3.18B and Qwen2-7B. Coverage for DRI is consistently higher and close to the reference summary compared to CANLII across all evaluated models. This pattern highlights the greater challenge of preserving salient argumentative information in legal texts, where arguments are often sparsely distributed across lengthy and complex contexts.
Error types We analyze the types of error extracted from of the ARC evaluation. Figure 3 shows that the most frequent error is missing, indicating that facts in arguments are often omitted rather than misrepresented. While factual inconsistencies exist, they are less prevalent than missing information. These findings emphasize that beyond hallucination, ensuring comprehensive salient information coverage remains a critical challenge in summarization.
5.2 Do argument positions in the source affect their coverage in summaries?
Positions of arguments in the source from the perspective of LLMs
@@ 5.2 Do argument positions in the source affect their coverage in summaries?
Following Ravaut et al. (2024b), we start by analyzing the positions where included LLMs look at in its context window. We leverage the lexical greedy approach for source sentences identification Ravaut et al. (2024a); Adams et al. (2023) by iteratively adding sentences in the source that maximizes ROUGE-1 score until there is no further improvement. We analyze the source sentence indices by their argument role annotations.
Figure 4 reveals a distinct U-shaped context window across all models in both datasets, with the effect being particularly pronounced in CANLII. Analyzing source sentences based on their argument role annotations suggests that argument positions are strongly influenced by this U-shaped pattern. This is especially concerning for CANLII, where reference summaries indicate that arguments do not adhere to a fixed positional pattern. In contrast, reference summaries in DRI more closely align with the LLM context window distribution.
Effect of context window positional bias on coverage
@@
We analyze the positions of salient arguments using . For CANLII, we apply greedy sentence selection while restricting the selection to the annotated arguments in both the reference summary and the input document, thereby reducing computational cost and ensuring that only arguments in the reference summary are mapped to arguments in the input. For DRI, we directly select the positions of arguments in the input document with a relevance score . We follow Ravaut et al. (2024b) by computing the mean relevant position of salient arguments and measuring the Pearson correlation between this mean and .
| Model | CANLII | DRI |
| LlaMA-3.1-8B | -0.230 | 0.129 |
| Mistral-8B | -0.369 | -0.055 |
| Qwen2-7B | -0.301 | -0.223 |
Table 5 999Correlation with full ARC scores included in Appendix G. shows a significant negative correlation in CANLII (), indicating that LLM context windows impact argument coverage. In DRI, correlations are weaker and not significant, with LlaMA-3.1-8B even showing a slight positive trend. These patterns align with Figure 4: reference summaries in CANLII differ in distribution from generated ones, while in DRI, argument roles better reflect the U-shaped context window bias. This suggests that context-window positional bias can hinder argument coverage, especially when arguments are sparsely distributed.
5.3 Do LLMs cover argument roles disproportionally?
We propose a bias score . This score is grounded in the metric. We compute for each argument role by first calculating its score and normalizing it by the role’s prior frequency in the source document. This adjustment corrects for overrepresented roles, ensuring the bias score reflects true disparities in coverage. The final bias score for an argument role is defined as:
Here, denotes the frequency of argument role in source document , and is the total number of arguments in . To mitigate bias from argument length and position—especially in CANLII, where longer roles and mid-position arguments negatively affected coverage—we compute within groups allowing up to word-length variation. We also control for position by selecting CANLII articles in which at least of arguments appear within the first or last of the case.
Figure 5 101010Appendix H includes bias analysis without prior frequency normalization to avoid denominator inflation. shows that salient arguments are not covered equitably across roles. In CANLII, conclusions are consistently better covered than issues and reasons across all groups. While length control reduces some role-related bias, conclusion remain the most covered, even under both length- and position-controlled settings. In contrast, for DRI, own claims are less covered relative to other roles, but when compared to reference summaries, their coverage is comparable. This suggests the lower bias score arises from normalization, reflecting the overrepresentation of own claims in the source corpus rather than a true coverage gap. Finally, while there is a bias to human reference in case of CANLII , the gap in argument coverage is notably larger for LLM outputs compared to DRI.
@@ 6 Conclusion and Future Work
We introduced ARC, a novel evaluation framework for evaluating how well LLM-generated summaries preserve salient arguments. Our multi-level formulation—spanning full-set, role-level, and subatomic coverage—correlates more strongly with human judgments compared to standard lexical, semantic, and entailment-based metrics, with subatomic evaluation showing highest overall correlation with human judgments. Using our proposed metric, we identified two key limitations in summarizing long legal opinions. First is positional bias, where LLMs tend to overrepresent content at the beginning or end of the input context, affecting the coverage of sparsely distributed arguments. Second is role bias, where models disproportionately favor covering conclusions over other roles such as issues and reasons. Future work can build on this work by exploring incorporating explicit argument representations during training or during prompting LLMs for a more informative summaries in high stakes domains.
@@
While the ARC framework enables a comprehensive multi-level evaluation of argument coverage, several limitations remain. First, our evaluation is constrained by the lack of dedicated benchmarks explicitly designed for argument coverage. Existing datasets offer limited annotation granularity, particularly at the subatomic level, which restricts the reliability of fine-grained assessments. As a result, our decomposition into atomic argument units depends on LLM-based prompting and entailment filtering, both of which may introduce inaccuracies. Second, the relatively small size of the DRI dataset (40 documents) limits the generalizability of our findings in the scientific domain. While we were constrained by the available datasets, developing larger, rigorously annotated datasets is an important direction for future work, though it is beyond the scope of this study. Finally, due to computational constraints, our evaluation only includes models that fit within available memory resources. We encountered out-of-memory issues with several larger models. Future work should consider incorporating larger-scale models to determine whether argument coverage improves with increased model capacity or if the observed limitations persist.
@@ A.4 Rhetorical roles per relevance to summary likert score
To better understand the relationship between rhetorical structure and relevance to the summary, we compute the percentage of each rhetorical role across Likert-rated sentences in the DRI corpus. As shown in Figure 8, non-argumentative content dominates among sentences rated as totally irrelevant to the summary (Likert score 1). However, as the perceived relevance increases, argumentative content becomes more prominent, with Own Claim consistently emerging as the most frequent rhetorical role across all higher-quality categories. This trend highlights a clear shift toward structured argumentative writing in more relevant reflections.
Appendix B Likert scores based on human evaluation
Table 8 shows the Likert scale from to definitions.
| Rating | Explanation of the Generated Summary Rating Scale |
Appendix C Fact decomposition algorithm
@@
Appendix D Evaluation Prompts
@@
Evaluation prompts for ARC , ARC, and ARC are described in Table 11, 12, 13 respectively. In each prompt, we ask the LLM to first generate a rationale before assigning a scoring decision following standard evaluation with LLMs Liu et al. (2023a).
| Prompt Given to GPT4-o for Fullset Evaluation |
| Task: |
| Evaluate how well a given summary covers a provided set of arguments. Assign a score from 1 to 4 based on the extent of coverage, provide a clear explanation for your rating, and output the result in a specified JSON format. |
| Instructions: |
| • Read the provided arguments and summary carefully. • Rate the extent to which the arguments are covered by the summary using the scale described in Table 8. • Format your evaluation as a JSON object with: – "explanation": A concise explanation of your rating. – "rating": The assigned score (1 to 4). |
| Example Output Format: |
| "explanation": "Place your explanation here", |
| "rating": "Place your rating here" |
| Input: |
| • Arguments: {reference_arguments} • Summary: {generated_summary} |
| Output: |
| Provide your evaluation in the specified JSON format. |
| Prompt Given to GPT4-o for Argument Role Evaluation |
| Task: |
| Determine whether a summary fully supports a given argument or omits/contradicts key information. |
| Instructions: |
| • Output 1 if the summary fully supports the argument without omissions or contradictions. • Output 0 if the summary fails to support the argument or contains contradictory or incorrect details (e.g., logical errors, entity mismatches). • Respond in a JSON object with: – "decision": Either 1 or 0. – "explanation": A brief justification, noting missing or conflicting content. |
| Input: |
| Argument: {argument} |
| Summary: {summary} |
| Output Format: |
| Respond only with a JSON object structured as: |
| "explanation": "<Brief reasoning for your decision>", |
| "decision": <0 or 1> |
| Note: Think critically before deciding. Do not include any extra text beyond the JSON output. |
@@
Appendix E Training and
We fine-tune DeBERTa, BigBirdRoBERTa, and LegalBERT using checkpoints obtained from the HuggingFace library Wolf et al. (2019). All models are trained using -fold cross-validation based on case-summary pairs for each model forming a total of case-summary pairs, resulting in argument-summary pairs for training and atomic fact-summary pairs for training . To handle class imbalance, we leverage weighted cross-entropy loss, where class weights are proportional to the label distribution across the dataset. This is particularly important for , where the label is significantly underrepresented compared to others. All models are trained for epochs using the Adam optimizer with an initial learning rate of . The maximum sequence length is set to for most models due to encoder constraints. However, for BigBirdRoBERTa, we expand the maximum length to to accommodate longer summaries and maintain coverage.
| Model | F1 Macro Score () | F1 Macro Score () |
| DeBERTa-base | 0.7138 | 0.7936 |
| DeBERTa-large | 0.7202 | 0.7936 |
| LegalBERT | 0.6850 | 0.7812 |
| BigBirdRoBERTa | 0.6629 | 0.7891 |
Table 14 shows that across all models, DeBERTa-Large obtained the best F1 macro average scores. Including encoders that can handle longer lengths didn’t improve the performance of prediction.
Appendix F ARC and ARC results
Figure 9 presents the results for both ARC and ARC across the CANLII and DRI datasets.
The ARC results indicate that both LlaMA-3.18B and Qwen2-7B achieve comparable levels of argument coverage from a holistic perspective. In contrast, and consistent with the ARC results presented in Section 5, Mistral-8B demonstrates slightly lower performance than both models. However, a key limitation of ARC is its insensitivity to fine-grained differences, as it primarily relies on coarse Likert-style scores that obscure nuanced variation across model outputs.
By contrast, ARC reveals that LlaMA-3.18B exhibits a higher number of perfectly covered roles compared to both Mistral-8B and Qwen2-7B. Nonetheless, due to its stricter definition of completeness at the role level, ARC underrepresents the strength of Qwen2-7B in capturing a greater number of atomic facts than Mistral-8B, as evidenced by both ARC and ARC.
Appendix G Correlation with source argument positions across full ARC scores
| Metric | Model | CANLII | DRI |
| ARC | LlaMA-3.1-8B | -0.137 | -0.152 |
| Mistral-8B | -0.042 | 0.250 | |
| Qwen2-7B | -0.187 | -0.174 | |
| ARC | LlaMA-3.1-8B | -0.200 | 0.202 |
| Mistral-8B | -0.244 | -0.065 | |
| Qwen2-7B | -0.216 | -0.124 | |
| ARC | LlaMA-3.1-8B | -0.230 | 0.323 |
| Mistral-8B | -0.369 | -0.055 | |
| Qwen2-7B | -0.301 | 0.041 |
Table 15 confirms that the LLM context window significantly impacts coverage across all scores (including ARC and ARC) in CANLII (), with a predominantly negative correlation. In contrast, its effect in DRI is not statistically significant, and in some cases (e.g., LlaMA-3.1-8B), it is slightly positive.

### v2

@@ 5.1 Obtaining Generated Summaries
We evaluate eight long-context open-weight LLMs from the LLaMA Grattafiori et al. (2024), Mistral Jiang et al. (2024) and Qwen Yang et al. (2024a); Team (2024) families that meet our computational constraints. Specifically, we include LLaMA-3.1-8B-Instruct, LLaMA-3.2-1B and 3B, Qwen-2.5-Instruct-1.5B, 3B, 7B, 14B, and Mistral-8B. We vary model sizes to examine whether scaling improves coverage performance. Inference is conducted using VLLM Kwon et al. (2023) for efficiency and scalability, with Qwen models further extended to tokens via RoPE scaling (factor ).
Summaries are generated using greedy decoding (), capped at tokens to ensure fair long-form generation. Each document is prompted with the following instruction: "Read the following text and summarize it: {input document}. Summarize in {reference summary word length} words. Summary:" We deliberately use a generic summarization prompt to evaluate LLMs without giving them any prior indication of which information is salient.777See Appendix G for experiments with an argument-aware prompt. This setup mirrors the human annotation process, where experts also lacked explicit guidance on saliency when producing reference summaries.
| CANLII100 | CANLII | DRI | ||||||||
| Model | MF | FE | MF | FE | MF | FE | MF | FE | MF | FE |
| Qwen-2.5-14B | 35.7 | 2.0 | 31.4 | 3.2 | 28.7 | 3.1 | 36.5 | 0.2 | 36.7 | 0.5 |
| Qwen-2.5-7B | 35.5 | 2.0 | 31.2 | 3.6 | 27.8 | 3.2 | 33.5 | 0.5 | 34.7 | 1.0 |
| Qwen-2.5-3B | 36.2 | 2.8 | 30.5 | 3.4 | 29.5 | 3.6 | 36.6 | 0.8 | 36.6 | 0.7 |
| Qwen-2.5-1.5B | 56.3 | 3.6 | 49.1 | 5.5 | 46.9 | 4.9 | 49.9 | 0.4 | 47.8 | 1.1 |
| Mistral-8B | 46.8 | 3.0 | 42.1 | 3.1 | 40.4 | 3.7 | 43.3 | 0.4 | 44.0 | 1.3 |
| LLaMA-3.1-8B | 40.1 | 2.1 | 36.9 | 3.6 | 33.1 | 3.3 | 41.1 | 0.3 | 41.7 | 0.8 |
| LLaMA-3.2-3B | 45.4 | 3.0 | 40.1 | 4.0 | 36.5 | 4.2 | 43.8 | 0.3 | 43.3 | 1.3 |
| LLaMA-3.2-1B | 56.3 | 4.6 | 49.1 | 5.5 | 45.1 | 5.7 | 49.9 | 0.7 | 50.0 | 1.3 |
The target length is dynamically matched to the reference summary for comparability. For DRI, the target length is fixed to the longest reference summary to encourage maximal argument retention.
5.2 RQ1: Can LLM summaries cover salient arguments without explicit supervision?
We compute ARCscore for all model-generated summaries described in Section 5.1, as well as for human reference summaries on both the CANLII and DRI benchmarks. For CANLII, we report results on both the smaller -article subset (CANLII100) scored with GPT-4o and the full set (CANLII) scored with DeepSeek-R1-Distill-Qwen14B, thereby testing the robustness of ARCscore across different judges and verifying that it assigns near-perfect scores to expert-written summaries. For DRI, because role-level annotations in summaries are unavailable, the goal is to assess how well human references capture salient argument roles compared to LLM-generated outputs.
Table 5 shows that argument coverage remains incomplete across all models and domains, indicating that LLMs could still benefit from argument-aware supervision. Although the two judges yield slightly different ARCscore values, the overall ranking of models is consistent. 888For CANLII100, Kendall’s (between the judges scores); for DRI, ; both . Two clear patterns emerge from the results. Larger models generally achieve higher coverage. Within the Qwen-2.5 family, performance improves when scaling from B to B and B across both judges. A similar trend holds for LLaMA-3.2, where the B model surpasses the B variant. 999This scaling effect is examined only for models with B–B parameters. Coverage is more challenging in legal opinions. On DRI, the strongest model (Qwen-2.5-7B) achieves , closely approaching the human reference () under both judges. By contrast, on CANLII100 and the full CANLII, the best-performing models plateau at and , far below the human reference ( and , respectively). This disparity underscores the difficulty of preserving salient argumentative content in legal texts, where arguments (both salient and non-salient) are sparsely distributed across lengthy contexts.
Fact-level Analysis. We analyze fact-level decisions (missing versus non-factual) to determine whether coverage errors arise primarily from omissions of key information or from factual inaccuracies. For each model, we compute the proportion in of missing facts and factual errors across all generated summaries. As shown in Table 6, missing facts are the dominant source of error across both datasets, judges, and models, indicating that salient information is more often omitted than misrepresented. While factual inconsistencies do occur, they are consistently less frequent. These findings suggest that, beyond hallucination, the central challenge in summarization is achieving comprehensive coverage of salient content.
5.3 RQ2: Do argument positions in the source affect their coverage in summaries?
@@
Following Ravaut et al. (2024), we start by analyzing the positions where included LLMs look at in its context window. We leverage the lexical greedy approach for source sentences identification Ravaut et al. (2024); Adams et al. (2023) by iteratively adding sentences in the source that maximizes ROUGE-1 score until there is no further improvement. 101010Appendix H contains the full algorithm. We analyze the source sentence indices by their argument role annotations. Figure 2 shows a clear U-shaped context window bias across all models, most pronounced in CANLII. Argument role analysis indicates that source positions are heavily shaped by this pattern—problematic for CANLII, where reference summaries do not follow fixed positional trends. In contrast, DRI references align more closely with the LLM bias distribution.
| CANLII100 | CANLII | DRI | |||
| Model | |||||
| Qwen-2.5-14B | -0.144 | -0.131 | -0.146 | 0.119 | -0.044 |
| Qwen-2.5-7B | -0.301 | -0.237 | -0.214 | -0.223 | 0.007 |
| Qwen-2.5-3B | -0.232 | -0.112 | -0.136 | 0.006 | 0.039 |
| Qwen-2.5-1.5B | -0.230 | -0.170 | -0.163 | 0.074 | -0.010 |
| Mistral-8B | -0.369 | -0.244 | -0.193 | -0.055 | 0.012 |
| LLaMA-3.1-8B | -0.230 | -0.088 | -0.153 | 0.129 | 0.166 |
| LLaMA-3.2-3B | -0.216 | -0.113 | -0.190 | -0.091 | -0.031 |
| LLaMA-3.2-1B | -0.171 | -0.087 | -0.198 | -0.020 | -0.031 |
@@
We analyze how the position of salient arguments affects coverage as measured by . For CANLII and CANLII100, we apply greedy sentence selection while restricting the pool to annotated arguments that appear in both the reference summary and the input document. This reduces computational cost and ensures that only arguments included in the reference summary are mapped back to the input. For DRI, we directly identify the positions of arguments in the input document with a relevance score of . Following Ravaut et al. (2024), we compute the mean relative position of salient arguments and measure the Pearson correlation between this mean position and .
Table 7 shows a consistent negative correlation in CANLII (both CANLII100 with GPT4-o and DeepSeek–R1, as well as the full CANLII set with DeepSeek–R1), with several models reaching significance (). This confirms that LLM context windows systematically bias coverage in long legal cases. By contrast, correlations in DRI are weaker and largely non-significant, with some models (e.g., LLaMA-3.1-8B) even showing slight positive trends under DeepSeek–R1. These results align with Figure 2: in CANLII, reference summaries diverge from the strong U-shaped bias present in model outputs, whereas in DRI, reference arguments more closely mirror the positional distribution induced by the LLM context window.
5.4 RQ3: Are argument roles disproportionately covered?
We propose as a role-specific bias score, designed so that higher values indicate stronger bias (i.e., poorer representation of a role). The intuition is that if a role is perfectly represented in the summary, then , and the bias should vanish (). Conversely, lower coverage implies a larger gap from perfect representation, reflected in a higher . Formally, we define:
where is the frequency of role in the source document , and is the total number of arguments in . The normalization term downweights roles that are overrepresented in the source, ensuring that captures true disparities in coverage rather than frequency effects.
To further reduce confounds from argument length and position—especially in CANLII,111111This analysis uses CANLII100 with the GPT4-o judge, as results in RQ1 and RQ2 show consistent trends across judges and dataset splits. where longer roles and arguments positioned in the middle of the document can bias coverage—we compute the bias score within length-controlled groups, allowing a variation in word count. In addition, to control for positional bias effect on coverage, we restrict the positional analysis to cases where at least 80% of arguments fall within either the first or the last 20% of the document. Positions are computed using the relative positions of the indices of argumentative sentences from the beginning of the document. These constraints help ensure that the observed effects on role coverage are not confounded by argument length or positional bias.
Figure 3121212Appendix I reports results without frequency normalization to rule out denominator inflation effects. shows that, for CANLII100, is consistently lowest for conclusions, indicating stronger coverage of this role across both the all-argument analysis and the position-controlled setting. Controlling for role length explains part of the variation across roles, but the relative advantage of conclusions persists. In DRI, by contrast, bias scores are generally lower than in CANLII100, consistent with our earlier finding that LLMs capture more salient arguments in scientific summaries. Nonetheless, background claims tend to be covered less effectively, even when length is controlled. Across both domains, model size further influences role-level disparities: for instance, models with B parameters consistently exhibit the highest , reflecting weaker and less balanced argument coverage.
@@ 6 Conclusion and Future Work
We introduced ARC, a hierarchical evaluation framework for assessing how well LLM-generated summaries preserve salient argumentative content. ARC provides a principled and interpretable diagnostic tool for evaluating structured argument coverage in long-context summarization. Its hierarchical structure not only enables interpretability—revealing which roles are preserved and whether errors arise from omissions or factual inaccuracies—but also achieves higher correlation in holistic evaluation compared to lexical, semantic, entailment, and decomposition-based baselines. Applying ARC to legal and scientific domains uncovered two consistent limitations of current LLMs: positional bias, where the characteristic U-shaped context window negatively affects coverage, and role bias, where conclusions are favored over other roles such as issues and reasons. Future work can extend this framework by incorporating explicit argument structures into training or prompting, and by leveraging ARC’s interpretable outputs to guide targeted improvements in summarization—particularly in high-stakes domains.
@@
While the ARC framework enables a comprehensive, multi-level evaluation of argument coverage, several limitations remain that suggest promising directions for future work.
Limited Benchmarks. Evaluation is constrained by the scarcity of benchmarks explicitly designed for argument coverage. Existing datasets provide limited annotation granularity, especially below the argument-unit level. To our knowledge, only the two benchmarks used here include both salience annotations and explicit argument roles. The small size of DRI (40 documents) also limits generalizability, motivating larger, rigorously annotated corpora—e.g., debates or financial texts—where arguments are central.
Decomposition Sensitivity. Our decomposition into atomic facts relies on the Yang et al. (2024b) algorithm, LLM-based prompting, and entailment filtering, which may introduce errors. While this decomposition proved more effective than that in FactScore Min et al. (2023), it is worth noting that FactScore relied on human-authored examples from biographical texts, which might have contributed to their lower correlations. Future extensions of ARC should consider incorporating expert-written atomic facts to further improve consistency and balance.
Balancing Precision and Recall. ARC currently emphasizes recall—whether salient argument-role facts are preserved—similar to FactScore. While critical for high-stakes domains, this ignores precision, i.e., over-inclusion of non-salient content. Future work should jointly assess both dimensions, for instance via a harmonic mean, for a fuller view of coverage quality. 131313We additionally conduct a precision–recall analysis on a subset of CANLII summaries; details and results are reported in Appendix J.
Dependence on Gold Annotations. Our analysis assumes access to gold salience labels. Though this allows conclusive evaluation, future research should explore end-to-end systems that integrate salience detection, fact decomposition, and coverage assessment within a unified framework.
@@
Acknowledgements
This material is based upon work supported by the National Science Foundation under Grant No. 2040490 and by Amazon. This research was supported in part by the University of Pittsburgh Center for Research Computing through the resources provided. This research was partially supported by the CS50 fellowship awarded by the department of computer science at the University of Pittsburgh. We want to thank the members of the Pitt PETAL group, Pitt NLP group, and anonymous reviewers for their valuable comments in improving this work.
@@ A.4 Rhetorical roles per relevance to summary Likert score
To better understand the relationship between rhetorical structure and relevance to the summary, we compute the percentage of each rhetorical role across Likert-rated sentences in the DRI corpus. As shown in Figure 6, non-argumentative content dominates among sentences rated as totally irrelevant to the summary (Likert score 1). However, as the perceived relevance increases, argumentative content becomes more prominent, with Own Claim consistently emerging as the most frequent rhetorical role across all higher-quality categories. This trend highlights a clear shift toward structured argumentative writing in more relevant argument roles.
Appendix B Fact decomposition algorithm
@@
Appendix C Atomic fact error examples
Table 12 presents representative examples from the DRI dataset, illustrating model decisions on atomic facts as evaluated by GPT-4o.
| Example 1 – Non-factual |
| Summary: The paper presents an efficient fluid simulation method using Laplacian eigenfunctions as a velocity basis, which avoids the high computational complexity of traditional Eulerian and Lagrangian approaches. This method is particularly advantageous for incompressible fluid dynamics, offering a principled way to reduce the dimensionality of the Navier–Stokes equations. The authors argue that Laplacian eigenfunctions provide a natural correspondence with spatial scales of vorticity, allowing precise control over turbulent energy at different scales. They show that for simple domains, these eigenfunctions have closed-form expressions, enabling fast and exact evaluation of the velocity field. The method is formulated using Galerkin projection onto Laplacian eigenfunctions, which are divergence-free and satisfy free-slip boundary conditions. For complex domains, a discrete exterior calculus (DEC) approach is used to compute the basis fields. The paper highlights the benefits of this method, including its ability to simulate fluid motion interactively, its low memory and precomputation requirements, and its flexibility in controlling viscosity and external forces. However, it also acknowledges limitations, such as stair-step artifacts on irregular boundaries and the need for additional methods to handle moving obstacles. The authors demonstrate that the method can be applied to various domains, including simple rectangles and complex 3D cavities, and discuss its potential for expressive control and optimization in fluid dynamics. The paper concludes by noting that while the method is particularly suited for gaseous phenomena, it may require further development for liquid simulations. |
| Fact: Our basis has a natural correspondence with spatial scales of vorticity that is lacking in [Treuille et al. 2006]. Decision: Non-factual Explanation: The summary correctly states that the basis corresponds to vorticity scales but omits the comparative qualifier “lacking in [Treuille et al. 2006],” misleadingly implying a novel rather than comparative contribution. |
| Example 2 – Missing Fact |
| Summary: Pose space deformation (PSD) is a technique that unifies and improves upon both shape interpolation and skeleton-driven deformation methods. It represents deformations as mappings from a pose space to displacements in the object’s local coordinate frames. PSD allows direct manipulation of desired shapes and maintains the performance benefits of traditional methods. Key applications include facial and body deformation in entertainment, telepresence, and gaming. The limitations of shape interpolation and skeleton subspace deformation (SSD) motivate the development of PSD. Shape interpolation lacks smoothness and requires frequent sculpting of new key shapes. SSD, while versatile, suffers from the "collapsing elbow" problem and is difficult to control. PSD addresses these issues by interpolating deformations in a pose space, enabling smooth and controllable deformations. PSD requires scattered data interpolation in high-dimensional spaces. Gaussian radial basis functions are used due to their well-behaved nature and ease of implementation. The algorithm is simple, general, and allows real-time synthesis, making it suitable for high-resolution models. Facial animation can benefit from PSD by allowing sculpting of intermediate expressions and multidimensional pose spaces. PSD can handle various deformation scenarios uniformly, from simple joints to complex secondary animations. The setup and synthesis costs are minimal, ensuring real-time performance. While anatomically based models offer higher quality, PSD provides a practical alternative for real-time applications and fanciful creature designs. |
| Fact: The synthesis cost is only slightly more than that of shape interpolation. Decision: Missing Fact Explanation: The summary omits the comparative detail that PSD’s synthesis cost is slightly higher than that of shape interpolation, overstating the efficiency of the proposed method. |
Appendix D Evaluation prompt
@@
Evaluation prompt for fact-level coverage ARC is described in Table 13. We ask the LLM to generate a rationale before assigning a decision.
@@
Appendix E Likert scores based on human evaluation
Table 14 shows the Likert scale from to definitions.
| Rating | Explanation of the Generated Summary Rating Scale |
Appendix F Full expert correlation
Table 15 reports correlations with individual experts and their averaged ratings across both judges. ARCscore (with DeepSeek-R1-Distill-Qwen14B) achieves the strongest overall correlation with expert judgments, surpassing all baselines. ARCscore (with GPT4-o) yields the second-highest overall correlation, outperforming all metrics except SummaCZS (sentence-level) on Expert 2. These results suggest that the decomposition in ARCscore not only enables fine-grained and interpretable analyses across error types and argument-role distinctions, but also preserves—and even improves upon—holistic coverage evaluation compared to strong automatic baselines.
| Metric | Expert 1 | Expert 2 | Average | |||
| ROUGE-1 | 0.378 | 0.520 | 0.347 | 0.456 | 0.391 | 0.539 |
| ROUGE-2 | 0.363 | 0.441 | 0.276 | 0.419 | 0.336 | 0.475 |
| ROUGE-L | 0.310 | 0.417 | 0.337 | 0.450 | 0.345 | 0.479 |
| BERTScore | 0.319 | 0.409 | 0.292 | 0.398 | 0.354 | 0.517 |
| (sent) | 0.310 | 0.517 | 0.427 | 0.517 | 0.387 | 0.537 |
| (doc) | 0.476 | 0.531 | 0.262 | 0.344 | 0.375 | 0.476 |
| (sent) | 0.329 | 0.371 | 0.341 | 0.460 | 0.345 | 0.459 |
| (doc) | 0.408 | 0.300 | 0.269 | 0.264 | 0.352 | 0.311 |
| FactScore (GPT4-o) | 0.365 | 0.511 | 0.362 | 0.460 | 0.405 | 0.549 |
| ARCscore (GPT-4o) | 0.499 | 0.607 | 0.401 | 0.482 | 0.465 | 0.593 |
| ARCscore (DeepSeek-R1) | 0.545 | 0.653 | 0.441 | 0.519 | 0.509 | 0.638 |
Appendix G Zero-shot summaries with argument-aware prompts
We perform a zero-shot ablation using four representative models from our set of eight: Llama-3.2-1B, Llama-3.2-3B, Qwen-2.5-3B, and Qwen-2.5-7B, varying both model families and sizes. In this setting, models are instructed to generate summaries of the same length as the human references, with an explicit emphasis on the key arguments in the input text. This experiment evaluates whether explicitly encoding argument salience in the prompt improves the inclusion of salient arguments in model-generated summaries. The zeroshot prompt used is described in Table 16.
| Argument-Aware Zero-Shot Prompt |
| Input: |
| Read the following text carefully: |
| {input} |
| Task Description: |
| Write an abstractive summary in about {length} words. |
| Instructions: |
| • Focus primarily on the key arguments presented in the text. • Ensure the summary is coherent, fluent, and written in continuous prose (not bullet points or key phrases). |
| Output: |
| Summary: |
Given the consistency across judges and datasets we report results on the smaller CANLII set (CANLII100) and the full DRI set using GPT-4-o as a judge.
Table 17 shows that explicitly instructing models to focus on key arguments did not consistently improve performance and, in some cases, even degraded coverage of salient roles. Only Qwen-2.5-3B,7B achieved a modest improvement on CANLII. These findings suggest that simply prompting models to emphasize arguments is insufficient for enhancing saliency coverage, underscoring the need for deeper alignment or fine-tuning strategies that more effectively encode argumentative relevance.
| Model | CANLII | DRI | ||
| Vanilla | Arg-Prompt | Vanilla | Arg-Prompt | |
| LLaMA-3.2-1B | 0.451 | 0.426 | 0.674 | 0.645 |
| LLaMA-3.2-3B | 0.573 | 0.570 | 0.705 | 0.667 |
| Qwen-2.5-3B | 0.648 | 0.654 | 0.730 | 0.718 |
| Qwen-2.5-7B | 0.676 | 0.703 | 0.799 | 0.754 |
Appendix H Source sentences identification
Algorithm 2 presents the lexical greedy procedure adopted from Ravaut et al. (2024) for identifying source sentences given a generated summary. We extend the original algorithm to return both the selected sentence indices and their corresponding argument role types, based on previously annotated sentence-level roles.